# Quality-Driven Curation of Remote Sensing Vision-Language Data via Learned Scoring Models

### v1

@@ 8.1 ScoreRS Training
Table 9: Hyperparameter for training our ScoreRS
We implement a three-stage training approach for our ScoreRS model. Throughout all stages, we employ the AdamW optimizer with cosine learning rate decay and set $(\beta_{1},\beta_{2})$ to $(0.9,0.95)$. For computational efficiency, we implement ZeRO-2 optimization with bfloat16 precision across all training stages. Additional hyperparameters are detailed in Tab. 9.
Figure 5: Representative examples from our image-caption preference dataset
8.2 CLIP Finetuning
@@
Evaluation Setting
We evaluate the fine-tuned CLIP model on both RS classification and RS image-text retrieval tasks. For the classification tasks, we employ text prompts in the format of ”a satellite photo of {class name}” and ”a satellite image of {class name}” to perform the classification.
8.3 Large VLMs Finetuning
@@
Stage-specific hyperparameters were as follows:
@@
We adhere to the evaluation dataset split established in the VHM and utilize their task-specific prompts for each evaluation task. To ensure fair comparison, we augment each prompt with the appropriate task identifier when evaluating models that support task identifier. For response generation across all models and tasks, we maintain consistent hyperparameters with a temperature of 1.0 and top-p of 1.0.
8.4 GRPO Training
For open-ended vision question answering tasks, we employ a binary accuracy reward: 1 if the prediction matched the ground truth, and 0 otherwise. For bounding box prediction tasks, we implement a graduated reward system based on IoU: 0.5 for IoU $>$ 0.5, 0.6 for IoU $>$ 0.6, 0.7 for IoU $>$ 0.7, and 1.0 for IoU $>$ 0.8, with 0 reward otherwise. The application of ScoreRS for close-ended questions has been previously described in Section 2.2, and the hyperparameter $\beta$ in Equation 3 is set to 0.2.
We employ LoRA on all linear layers of the base LLM with a rank of 128 and $\alpha$ of 128. During the generation phase of GRPO, we configure the sampling temperature and top-p parameter to 0.9, with a maximum token generation limit of 256. We set the KL penalty coefficient to 0.04 and use a batch size of 8, with each sample generating 5 candidate answers for reward calculation.
We optimize the model using AdamW with a cosine learning rate schedule and a base learning rate of $8\times 10^{-6}$. To improve computational efficiency, we utilized bfloat16 precision, gradient accumulation with a factor of 4, and a maximum image resolution of 1024 pixels. Memory consumption is reduced by implementing Flash Attention in conjunction with ZeRO-2 optimization. All experiments are conducted for a single epoch.
For the supervised finetuning of Qwen2VL-7B-RS with our manually collected reasoning data, we employ identical hyperparameter settings as described in Sec. 8.3, with the LoRA rank set to 128.
Figure 6: Representative examples from our vision instruction preference dataset
Evaluation Setting
@@ 8.4 GRPO Training
The evaluation metrics, datasets, and framework for assessing the trained reasoning model remained consistent with those detailed in Sec. 8.3. The key distinction is the implementation of an answer extraction process from the reasoning model’s output for evaluation purpose. Importantly, if this parsing process failed to extract a valid answer from a reasoning model, we classify the response as incorrect.
9 Qualitative Example
@@
We provide representative conversation examples generated by our finetuned model, Qwen2VL-7B-RS-R1, in Fig. 7.
10 Reasoning for Reference Based Reward Calculation
@@
In Sec. 2.2, we introduce the reference-based reward mechanism using our ScoreRS for utilizing open-ended questions. Before arriving at this approach, we explored multiple strategies. The following describes our unsuccessful attempts and our reasoning:
@@
We initially attempted to directly use the output of ScoreRS for rewarding. However, since the output of ScoreRS can be significantly larger or smaller than the 0-1 range, and we do not know the possible range of the ScoreRS output (therefore, could not normalize them), the training proved very unstable. Additionally, the rule-based reward for close-ended questions was greatly overshadowed.
@@
We then tried normalizing the ScoreRS output with sigmoid or other similar normalization functions (tanh). This approach was problematic because the range of our ScoreRS output likely falls in the convergence part of these functions (for example, in the $>$ 10 or $<$ -10 regions of the sigmoid function). Although we found ways to sketch the domain of the function, learning remained too slow and did not yield satisfying results.
@@
With these unsuccessful attempts behind us, we opted for the reference-based reward calculation. Although the requirement for reference datasets may reduce the possible data size for RL training, we conclude that data volume is not the critical part of RL training—better answer sampling and reward strategies are far more important. Of course, we do not consider this the optimal solution, but rather our current finding, and we hope future work will explore more effective approaches.
11 Limitation and Discussion
@@
In this work, we demonstrate the efficacy of parameter-wise learned scoring models for RS vision-language data selection. Our investigation reveals that current RS vision-language datasets fall considerably short of optimal quality, underscoring the need for more rigorous curation efforts in this domain-specific context.
@@
As the machine learning community increasingly validates the ”less is more” principle [15, 53, 45], the RS community should prioritize data quality improvement through systematic quality control mechanisms or scoring model implementations. While our work highlights the suboptimal quality of existing RS vision-language data, we have not exhaustively explored what constitutes ”better” or ”more valuable” data for finetuning VLMs in the RS domain. Future research should investigate various dimensions including data difficulty levels, category distributions, and optimal data combinations to fully harness the potential of RS-specialized VLMs.
@@
Throughout our investigation, we gained significant insights regarding the misalignment between current data characteristics and the capabilities expected from advanced VLMs. Modern large-scale VLMs excel not only in perception tasks but also in reasoning and planning for complex problem-solving [52]. For the RS community, these models should ideally support sophisticated applications such as disaster analysis, urban planning, and transportation system assessment. However, current datasets predominantly focus on basic perception capabilities—classification, object localization, and color detection—tasks that could potentially be addressed by more lightweight models.
@@
System Message: You are a highly precise and analytical image description evaluator. Your task is to select the most accurate caption from given options or provide your own if none are satisfactory. You should approach this task systematically and provide detailed justification while maintaining objectivity. Prompt: Please analyze the given image and evaluate these three captions: Caption 1: {caption1} Caption 2: {caption2} Caption 3: {caption3} Follow these steps: 1. First, provide your own detailed description of the image, focusing on: - Observable objects and their characteristics - Spatial relationships and positioning - Key visual patterns and structures - Colors and textures Avoid making assumptions about image type or source. 2. Evaluate each provided caption using the following criteria (score 1-10): - Accuracy: How well does it match the visible elements? - Completeness: Does it capture all significant features? - Conciseness: Is it appropriately detailed without unnecessary information? - Objectivity: Does it avoid assumptions or hallucinated details? - Spatial Clarity: How well does it describe object relationships and positions? 3. Provide your analysis in the following format: [description] your independent description [end-description] [evaluation] 1: [score],[score],[score],[score],[score],[justification] 2: [score],[score],[score],[score],[score],[justification] 3: [score],[score],[score],[score],[score],[justification] [end-evaluation] [selection] best-id: [id] best-reason: [reason] [end-selection]
@@
Prompt: You are an expert evaluator specialized in determining the quality of answers. Your task is to systematically analyze and compare answers to determine which one is better. You must: 1. Always analyze the core requirements of the question first 2. Break down both the given answers into key components 3. Compare them systematically and objectively 4. Provide conclusions in a strictly formatted output and output your confidence Your evaluation must be based on: - Completeness: All key points covered - Accuracy: Correctness of information - Relevance: Direct address of the question - Clarity: Clear communication of ideas - Hallucination: Whether the answer is hallucinated Now, please evaluate the following answer pair: QUESTION:{question} ANSWER: {answer1} {answer2} {answer3} Provide your evaluation with the following structure and the result as json object: “‘json { ”answer1”: { ”Completeness”: [0-5] ”Accuracy”: [0-5] ”Relevance”: [0-5] ”Clarity”: [0-5] ”Hallucination”: [0-5] }, ”answer2”: … ”answer3”: … ”confidence”: [0-5] } “‘
@@
You are an expert in remote sensing and satellite image analysis. Transform the following question into a new, more diverse version. The new question should: 1. Test the same core concept but approach it differently 2. Either increase or decrease the complexity 3. Change the context or application domain 4. Use a different response format or analytical approach Also please task care that there is just single image. So do not output any temporal question. Moreover, do not explicitly mention the image is satellite image. Original question: ”{question}” Generate ONE new question that maintains the spirit of the original but differs in at least 2 of the above aspects. Do not explain your choices - only output the new question. Output here:

### v2

@@ D.2 ScoreRS Training
We implement a three-stage training approach for our ScoreRSmodel. Throughout all stages, we employ the AdamW optimizer with cosine learning rate decay and set $(\beta_{1},\beta_{2})$ to $(0.9,0.95)$. For computational efficiency, we implement ZeRO-2 optimization with bfloat16 precision across all training stages. Additional hyperparameters are detailed in Table 15.
D.3 CLIP Finetuning
@@
D.3.1 Evaluation Setting
We evaluate the fine-tuned CLIP model on both RS classification and RS image-text retrieval tasks. For the classification tasks, we employ text prompts in the format of "a satellite photo of {class name}" and "a satellite image of {class name}" to perform the classification.
Evaluation Metric:
For the classification task, we report top-1 and top-5 accuracies. For the retrieval task, we evaluate both image-to-text and text-to-image performance using top-1 and top-5 recall metrics.
Evaluation Dataset:
(1) NWPU (NWPU-RESISC45) [9] contains 31,500 RGB images devided into 45 scene classes, each class containing 700 images. We use the entire dataset for evaluation. (2) EuroSAT [20] is based on Sentinel-2 satellite imagtes consisting out of 10 classes with in total 27,000 labeled images. We use the entire dataset for evaluation. (3) fMoW [10] contains 1,047,691 images convering 207 country and the total images are devided into 63 categories. We use the test split of fMoW-RGB for evaluation (16,948 samples). (4) AID [67] contains 10,000 images from 30 different categories. We use the entire dataset for evaluation. (5) SIRI-WHU [37] is a classfication dataset that contains 2,400 images from 12 classes, with each class has 200 images. We use the entire dataset for evaluation. (6) WHU-RS19 [13] contains 19 classes, with each class has about 50 samples. We use entire dataset for evaluation. (7) UCM (UCM-Caption) [34] contains 2,100 images, with each image has 5 different captions. We use the same retrieval evaluation recipe as Skyscript [62]. (8) RSICD [34] contains 10,921 images, with each image has 5 different captions. We use the same retrieval evaluation recipe as Skyscript [34].
Figure 10: Representative examples from our image-caption preference dataset. Green represents reasonably good expression, while red represents low-quality expression
Figure 11: Representative examples from our vision instruction preference dataset. Green represents reasonably good expression, while red represents low-quality expression
D.4 Large VLMs Finetuning
@@
Stage-specific hyperparameters are as follows:
@@
D.4.1 Evaluation Setting
We adhere to the evaluation dataset split established in the VHM [44] and utilize their task-specific prompts for each evaluation task. To ensure fair comparison, we augment each prompt with the appropriate task identifier when evaluating models that support task identifier. For response generation across all models and tasks, we maintain consistent hyperparameters with a temperature of 1.0 and top-p of 1.0.
Evaluation Dataset:
(1) Classification: Brief introductions to each classification dataset can be found in Section D.3.1. (2) RSVQA[33]: RSVQA is a RS vision question answering dataset available in high resolution (RSVQA-HR) and low resolution (RSVQA-LR) versions. RSVQA-HR contains four question types: comparison (comparing object quantities), presence (determining if images contain specific objects), counting (quantifying objects), and area (measuring object areas). RSVQA-LR includes an additional rural/urban type to classify image settings. Following established practices in existing RS-specific VLMs such as VHM[44], SkysenseGPT [36], and GeoChat [27], we excluded questions related to counting and area measurement for evaluation. (3) Grounding: We used VG-DIOR [74] for vision grounding tasks. VG-DIOR is a RS vision grounding dataset built from RSVG [74] on the DIOR RS detection dataset [28]. Each sample contains an object description that requires the model to predict object coordinates. We calculated the IoU between predicted and ground truth coordinates, considering predictions with IoU greater than 0.5 as correct. (4) General Knowledge: We used LHRS-Bench [39] to holistically evaluate VLMs across different domains. LHRS-Bench is a multiple-choice question answering dataset covering 11 evaluation dimensions (from basic recognition to complex reasoning). For evaluation, we used the following prompt template: Question. Please answer the above question with the given choice (just answer with choice index): Choices. We extracted choice indices (e.g., A, B, C, D) using regular expressions and only considered responses containing the exact choice index as correct. Responses that included the content of the choices were not counted as correct. We did not implement the circular evaluation protocol introduced in [39].
D.5 GRPO Training
For open-ended vision question answering tasks, we employ a binary accuracy reward: 1 if the prediction matched the ground truth, and 0 otherwise. For bounding box prediction tasks, we implement a graduated reward system based on IoU: 0.5 for IoU >> 0.5, 0.6 for IoU >> 0.6, 0.7 for IoU >> 0.7, and 1.0 for IoU >> 0.8, with 0 reward otherwise. The application of ScoreRS for close-ended questions has been previously described in Section 3.3, and the hyperparameter $\beta$ in Equation 3 is set to 0.2. We also weight the format reward at 0.3 following the practice established in R1-VL [76].
We conduct full parameters training during the GPRO RL process. During the generation phase of GRPO, we configure the sampling temperature to 0.9, and top-p to 0.99. The maximum generated tokens are limited to 4096. We set the KL penalty coefficient to 0.01 and use a batch size of 16, with each sample generating 5 candidate answers for reward calculation.
We optimize the model using AdamW with a cosine learning rate schedule and a base learning rate of $1\times 10^{-6}$. To improve computational efficiency, we utilized bfloat16 precision, gradient accumulation with a factor of 2, and a maximum image resolution of 1024 pixels. Memory consumption is reduced by implementing Flash Attention in conjunction with FSDP strategy [80]. All experiments are conducted for 2 episodes.
For the supervised finetuning of Qwen2VL-7B-RS with our manually collected reasoning data, we employ identical hyperparameter settings as described in Section D.4, with the LoRA rank set to 128.
@@ D.5 GRPO Training
The evaluation metrics, datasets, and framework for assessing the trained reasoning model remained consistent with those detailed in Section D.4. The key distinction is the implementation of an answer extraction process from the reasoning model’s output for evaluation purpose. Importantly, if this parsing process failed to extract a valid answer from a reasoning model, we classify the response as incorrect.
Appendix E Qualitative Example
E.1 Ranked Data Comparison
In order to provide a qualitative comparison of samples receiving higher scores from different scoring models, we plotted the highest and lowest ranked samples from RemoteCLIP, VHM image-caption, and VHM vision instruction datasets in Figure 12.
From these results, we observe that for RemoteCLIP image descriptions, ScoreRS prefers more detailed descriptions while assigning lower scores to samples containing incorrect qualitative descriptions or claims that cannot be verified from the image. For the VHM image-caption dataset, ScoreRS can precisely identify accurate descriptions while filtering out incorrect or hallucinated content more effectively than CLIP-score.Finally, regarding the VHM vision instruction dataset, we clearly see that ScoreRS favors samples that demonstrate high-level reasoning and relevance to specific applications, while effectively filtering out incorrect answers.
E.2 Qwen2VL-7B-RS-R1 Inference Example
@@
We provide representative conversation examples generated by our finetuned model, Qwen2VL-7B-RS-R1, in Figure 13.
Appendix F Discussion nand Limitation
F.1 CLIP-Score or ScoreRS?
The results in Table 1, Table 2, and Table 10 indicate that the performance improvement achieved through data filtering with ScoreRS may be somewhat marginal compared to CLIP-score (approximately 1% improvement on RemoteCLIP data, while 4% on Skyscript). However, it is important to acknowledge that the RemoteCLIP and Skyscript datasets typically contain only short, brief descriptions (e.g., A ship in the middle of the picture), which aligns with the type of text captions CLIP is trained on [48, 69]. Furthermore, during evaluation on classification tasks, we typically employ simple text templates such as A photo of {class name} as the representative embedding for class matching. The descriptions used in retrieval evaluation tests are similarly short and concise. In contrast, the preference data used to train our ScoreRS typically consists of longer, more informative text (Figure 10), which may explain why ScoreRS does not show substantial advantages when qualifying these particular datasets. Nevertheless, even in these scenarios, our ScoreRS outperforms CLIP-score, demonstrating the significant potential of using VLMs as scoring models for data qualification.
We should acknowledge that for applications and datasets similar to RemoteCLIP or Skyscript, using CLIP-score is a reasonable approach. However, for qualifying more complex vision-language data and for advanced applications such as BoN selection and reinforcement learning, ScoreRS offers superior performance, as evidenced by Table 3, Figure 4, and Table 5.
F.2 Reasoning for Reference Based Reward Calculation
@@
In Section 3.3, we introduce the reference-based reward mechanism using our ScoreRS for utilizing open-ended questions. Before arriving at this approach, we explored multiple strategies. The following describes our unsuccessful attempts and our reasoning:
@@
We initially attempted to directly use the output of ScoreRS for rewarding. However, since the output of ScoreRS can be significantly larger or smaller than the 0-1 range, and we do not know the possible range of the ScoreRS output (therefore, could not normalize them), the training proved very unstable. Additionally, the rule-based reward for close-ended questions was greatly overshadowed. Although we acknowledge the potential challenges and attempt to mitigate them during the training process of our ScoreRS, the results presented in Appendix C.1 demonstrate that using raw scores for training provides the best training stability.
@@
We then try normalizing the ScoreRS output with sigmoid or other similar normalization functions (tanh). This approach was problematic because the range of our ScoreRS output likely falls in the convergence part of these functions (for example, in the >> 10 or << -10 regions of the sigmoid function). Although we find ways to sketch the domain of the function, learning remains too slow and does not yield satisfying results.
@@
With these unsuccessful attempts behind us, we opt for the reference-based reward calculation. Although the requirement for reference datasets may reduce the possible data size for RL training, we conclude that data volume is not the critical part of RL training—better answer sampling and reward strategies are far more important. Of course, we do not consider this the optimal solution, but rather our current finding, and we hope future work will explore more effective approaches.
F.3 RS Vision-Language Data Quality
@@
In this work, we have developed a framework to evaluate the quality of RS vision-language data and demonstrated the effectiveness of parameter-wise learned scoring models for RS vision-language data selection. Our investigation reveals that current RS vision-language datasets fall considerably short of optimal quality, underscoring the need for more rigorous curation efforts in this domain-specific context.
@@
As the machine learning community increasingly validates the "less is more" principle [25, 82, 72], the RS community should prioritize data quality improvement through systematic quality control mechanisms or scoring model implementations. While our work highlights the suboptimal quality of existing RS vision-language data using our ScoreRS, we acknowledge that ScoreRS is merely an initial attempt to qualify such data. Future research should investigate whether using separate scoring models for each quality dimension introduced in Appendix A could bring additional benefits. Further exploration should also address various dimensions including data difficulty levels, category distributions, and optimal data combinations to fully harness the potential of RS-specialized VLMs.
F.4 Expectation for Building RS-Specific VLMs
Throughout our investigation and evaluation, we gained significant insights regarding the misalignment between current data characteristics and the capabilities expected from advanced VLMs.
@@ F.4 Expectation for Building RS-Specific VLMs
Modern large-scale VLMs excel not only in perception tasks but also in reasoning and planning for complex problem-solving [81]. For the RS community, these models should ideally support sophisticated applications such as disaster analysis, urban planning, and transportation system assessment. Our analysis suggests that current RS-specific VLMs, while performing admirably on basic perception tasks, fall short on more complex challenges compared to general VLMs (Table 4).
Since basic perception tasks can be effectively addressed with more lightweight models (for example, accuracy on AID can achieve over 95% with the much smaller ViT-B architecture [60]), the development of large RS-specific VLMs could, in our view, put more focus on expanding beyond these foundational capabilities toward building agent-like assistants that interact with RS experts to conduct geological information analysis. This requires equipping models with more advanced capabilities such as planning, task decomposition, and reflection. Therefore, while continuing to value and build upon the important work in perception tasks, we suggest that future RS-specific VLM development could benefit from increased emphasis on these more complex reasoning capabilities to fully leverage the potential of LLMs in RS applications.
@@
Figure 12: Ranked examples from different dataset with different score models. Red represents low-quality expression
@@
Table 16: Scoring system for RS image-caption
Aspect Score Definition for Descriptions/Captions
1. Relevance 1 (Poor) The description is largely irrelevant to the image, contains significant hallucinations (objects/features clearly not present), or fundamentally misrepresents the scene.
2 (Weak) The description has notable inaccuracies, some hallucinated minor elements, or only vaguely relates to the primary content of the image.
3 (Fair) The description is generally relevant and mostly accurate, but may contain minor inaccuracies, slight misinterpretations, or omit verification of some mentioned details.
4 (Good) The description is accurate, clearly relevant, and faithfully represents the main visual content with no significant hallucinations or misinterpretations. Minor, non-critical details might be simplified.
5 (Excellent) The description is perfectly accurate, highly relevant, and precisely reflects all key visual elements and their relationships in the image without any ambiguity or hallucination.
2. Specificity & Detail 1 (Poor) The description is extremely vague (e.g., "an image of the ground," "some features") and provides almost no useful detail about the RS image content.
2 (Weak) The description is mostly general, lacking specific details about objects, patterns, or spatial arrangements that are clearly visible.
3 (Fair) The description provides some specific details but remains general in many aspects. Key distinguishing features might be mentioned but not elaborated upon.
4 (Good) The description offers a good level of specific detail about important objects, their characteristics (if discernible), and their general layout.
5 (Excellent) The description is rich in specific, fine-grained details about objects, textures, shapes, counts (where appropriate), and their precise spatial relationships.
3. Completeness 1 (Poor) The description misses almost all salient and important features/regions in the RS image.
2 (Weak) The description omits several significant features or covers only a small, non-representative part of the image.
3 (Fair) The description covers some of the main features but misses other notable ones or lacks comprehensive coverage of the overall scene.
4 (Good) The description covers most of the salient features and important regions of the image adequately.
5 (Excellent) The description provides a comprehensive account of all major salient features, land cover types, and significant patterns visible across the entire image.
4. Clarity & Fluency 1 (Poor) The description is largely incomprehensible, grammatically incorrect, or uses language so poorly that its meaning is lost.
2 (Weak) The description is difficult to understand due to significant grammatical errors, awkward phrasing, or unclear language.
3 (Fair) The description is mostly understandable but contains some grammatical errors, awkward phrasing, or minor ambiguities.
4 (Good) The description is clear, grammatically correct, and well-phrased, making it easy to understand.
5 (Excellent) The description is exceptionally clear, concise, grammatically flawless, and uses precise, fluent language.
5. Semantic Richness 1 (Poor) The description uses no relevant remote sensing terminology or misuses terms completely. Lacks any understanding of RS-specific concepts.
2 (Weak) The description uses very generic terms (e.g., "green areas," "buildings") with minimal or incorrect RS-specific vocabulary.
3 (Fair) The description uses some basic and appropriate RS terminology (e.g., "urban area," "farmland") but lacks depth or precision in describing RS-specific features.
4 (Good) The description correctly employs relevant RS terminology to describe features, land cover, or patterns (e.g., "center-pivot irrigation," "industrial zone").
5 (Excellent) The description expertly uses precise and advanced RS terminology, accurately identifying and describing complex features, phenomena, or sensor characteristics relevant to the image.
Table 17: Scoring system for RS instructions samples (Part 1)
Aspect Score Definition for Instruction/QA Pairs
1. Relevance 1 (Poor) The answer is completely irrelevant to the question, is based entirely on hallucinated information not in the image, or confidently answers an unanswerable question incorrectly.
2 (Weak) The answer poorly addresses the question, contains significant inaccuracies based on the image, or makes large, unfounded assumptions. Wrongly asserts unanswerable question is answerable.
3 (Fair) The answer attempts to address the question and is mostly based on the image, but may have minor inaccuracies, misinterpretations, or if the question is unanswerable, it may fail to state this clearly or attempt a guess.
4 (Good) The answer accurately and directly addresses the question using information verifiable in the image. If the question is unanswerable from the image, it states so clearly.
5 (Excellent) The answer perfectly and precisely addresses all aspects of the question using only visual evidence from the image. If unanswerable, it explicitly and correctly states why based on image content.
2. Specificity & Detail 1 (Poor) The answer is extremely vague (e.g., "Yes," "Maybe," "Objects are present") and provides no specific information from the image relevant to the question.
2 (Weak) The answer is too general and lacks specific details that are visible in the image and pertinent to answering the question effectively.
3 (Fair) The answer provides some relevant detail but could be more specific or elaborate further based on visible image content and the question’s needs.
4 (Good) The answer provides a good level of specific detail directly from the image that sufficiently addresses the question.
5 (Excellent) The answer is highly specific and rich in relevant details extracted from the image, providing a precise and thorough response to the question.
3. Completeness 1 (Poor) The answer completely fails to address the core of the question or ignores key components of a multi-part question.
2 (Weak) The answer addresses only a small part of the question or provides a very superficial response, missing obvious follow-up details implied by the question and visible in the image.
3 (Fair) The answer addresses the main part of the question but may be incomplete, missing some nuances, or not fully utilizing available visual information. If question is partially unanswerable, this might not be fully clarified.
4 (Good) The answer comprehensively addresses all explicit parts of the question using available image information. Clearly indicates if parts are unanswerable.
5 (Excellent) The answer fully and exhaustively addresses all aspects of the question, including implicit sub-questions where appropriate, based on thorough image interpretation. Clearly delineates what can and cannot be answered.
4. Clarity & Fluency 1 (Poor) The question is incomprehensible, or the answer is grammatically nonsensical, making the entire QA pair useless.
2 (Weak) The question is ambiguous, or the answer is poorly phrased with significant grammatical errors, making the QA pair difficult to understand or trust.
3 (Fair) The question is mostly clear, and the answer is generally understandable but may contain minor grammatical errors, awkward phrasing, or slight ambiguities.
4 (Good) The question is clear and unambiguous. The answer is well-phrased, grammatically correct, and easy to understand.
5 (Excellent) The question is exceptionally clear and well-posed. The answer is perfectly fluent, concise, grammatically flawless, and directly responsive.
Table 18: Scoring system for RS instructions samples (Part 2)
Aspect Score Definition for Instruction/QA Pairs
5. Semantic Richness 1 (Poor) Neither the question nor the answer uses any relevant RS terminology, or terms are severely misused, showing no domain understanding.
2 (Weak) The QA pair uses very generic terms. If RS terms are used, they are minimal, overly simplistic for the context, or slightly incorrect.
3 (Fair) The question or answer uses some basic, appropriate RS terminology, but could be more precise or leverage more domain-specific knowledge relevant to the image.
4 (Good) The question and/or answer correctly employ relevant RS terminology that enhances the specificity and technical accuracy of the exchange.
5 (Excellent) The QA pair expertly uses precise and potentially advanced RS terminology. The question might probe RS-specific insights, and the answer provides them accurately.
@@
System Message: You are a highly precise and analytical image description evaluator. Your task is to select the most accurate caption from given options or provide your own if none are satisfactory. You should approach this task systematically and provide detailed justification while maintaining objectivity. Prompt: Please analyze the given image and evaluate these three captions: Caption 1: {caption1} Caption 2: {caption2} Caption 3: {caption3} Follow these steps: 1. First, provide your own detailed description of the image, focusing on: - Observable objects and their characteristics - Spatial relationships and positioning - Key visual patterns and structures - Colors and textures Avoid making assumptions about image type or source. 2. Evaluate each provided caption using the following criteria: The ceiteria in Table 16. 3. Provide your analysis in the following format: [description] your independent description [end-description] [evaluation] 1: [score],[score],[score],[score],[score] 2: [score],[score],[score],[score],[score] 3: [score],[score],[score],[score],[score] [end-evaluation] [selection] best-id: [id] best-reason: [reason] [end-selection]
@@
Prompt: You are an expert evaluator specialized in determining the quality of answers. Your task is to systematically analyze and compare answers to determine which one is better. You must: 1. Always analyze the core requirements of the question first 2. Break down both the given answers into key components 3. Compare them systematically and objectively based on the given criteria 4. Provide conclusions in a strictly formatted output and Your evaluation must be based on the following criteria: The criteria in Table 17 and Table 18 Now, please evaluate the following answer pair: QUESTION:{question} ANSWER: {answer1} {answer2} {answer3} Provide your evaluation with the following format: [evaluation] 1: [score],[score],[score],[score],[score] 2: [score],[score],[score],[score],[score] 3: [score],[score],[score],[score],[score] [end-evaluation] [selection] best-id: [id] best-reason: [reason] [end-selection]
@@
You are an expert in remote sensing and satellite image analysis. Transform the following question into a new, more diverse version. The new question should: 1. Test the same core concept but approach it differently 2. Either increase or decrease the complexity 3. Change the context or application domain 4. Use a different response format or analytical approach Also please task care that there is just single image. So do not output any temporal question. Moreover, do not explicitly mention the image is satellite image. Original question: "{question}" Generate ONE new question that maintains the spirit of the original but differs in at least 2 of the above aspects. Do not explain your choices - only output the new question. Output here: